# Evidence Contextualization and Counterfactual Attribution for Conversational QA over Heterogeneous Data with RAG Systems

### v2

@@ 6. The ConfQuestions Benchmark
Need for a new benchmark. While there are many public QA benchmarks (Roy and Anand, 2022), there were none that fit all our desiderata: uses heterogeneous pages in enterprise wikis, suitable for ConvQA with RAG, and contains complex questions in more than one language. So we create our own benchmark ConfQuestions as follows. We crawled $10$ Spaces under https://openxt.atlassian.net/wiki/spaces, yielding $215$ public Confluence pages from the developer Atlassian. These contain several documents that are very close to each other content-wise (like test reports for software versions $6-9$, organizational policy over several years, several meeting notes on a topic) yet each page contains unique information: this is very suitable for building accurate retrieval and generation models. The pages contain a mix of concepts like software compatibility, build configurations, planning, along with mentions of people, dates and software, manifested via heterogeneous elements like tables, lists, and passages. These documents thus nicely simulate an enterprise Wiki space, our focus here.
Benchmark construction. The authors went through every page to judge the feasibility of generating NL conversations over them. Eventually, $50$ conversations were generated: $80\%$ with 5 turns, $20\%$ longer ones with $10$ turns, making context handling tougher, so $300$ questions in all. $20\%$ conversations also span two URLs instead of one, containing slight topic shifts. The annotators provided questions, intent-explicit versions, answer source (passage/list/table), complexity type (simple/complex), and ground truth answers along with source URLs. The questions were originally generated in English (EN). Native German speakers then translated both conversational and completed questions by hand to German (DE). We took special care to make the benchmark challenging from several aspects: (i) $50\%$ the questions are complex in the sense that they either require joining multiple evidences or involve aggregations, negations, or comparisons; (ii) the answers to these questions are spread equally across passages, lists, and tables. So to be able to perform well overall on ConfQuestions, a RAG systems needs to be able to handle complex questions over heterogeneous elements in multiple languages. Benchmark statistics are in Table 1.
@@ 7. Results and Insights
Data $\downarrow$ / Context $\rightarrow$ NONE +TTL +HDR +BEF +AFT +ALL NONE +TTL +HDR +BEF +AFT +ALL
@@ 7.1. Metrics
We use Precision@1 as our retrieval metric (since there is usually one gold URL), and Answer relevance as judged by gpt-4o (00 for non-relevant, $0.5$ for partially relevant, $1$ for relevant) against gold answers, as our metric for answer quality. Question-wise results are averaged over the full dataset. We use accuracy for evaluating attribution explanations as follows: we find the evidence that has the highest attribution score, and find its source URL. If this URL matches that of the gold answer, then accuracy is $1$, else 00. The average accuracy over all questions is then reported, that is also interpretable as a percentage if multiplied by $100$.
@@ 7.2. Setup
We are now ready to run and evaluate RAGonite in various configurations over the $600$ English and German questions in ConfQuestions. The proposed default configuration uses: (i) all available context; (ii) all evidences as the corpus; (iii) LLM-generated intent-explicit questions; (iv) verbalization as the linearization technique for table records; (v) both row- and table-level indexing for table elements; (vi) GPT-4o as our LLM of choice; (vii) BGE is as our text embedding model; (viii) hybrid search as our retrieval mode; and (ix) BGE reranking over the hybrid search results. Contextualization variants and attribution results are discussed in this section: all other ablation analysis is postponed to Sec. 7.4.
Prompts. There are three prompts used in RAGonite: (i) for rephrasing a conversational question; (ii) for generating an answer from retrieved evidences; and (iii) for evaluating a generated answer. These are presented below in Figs. 67, and 8, respectively (used for both English and German questions).
@@ 7.3. Key findings
Evidence contextualization improves performance. Contextualization results are in Table 2. We make the following observations: (i) Adding all the suggested contexts (ALL) consistently led to significantly improved performances for retrieval and generation. This means that enhancing evidences from documents with surrounding context is generally beneficial for helpful for RAG over collaborative wikis: content authors are often efficiency-oriented and do not repeat information inside passages, lists, or tables that are already mentioned in page titles or previous headings, for example. Similarly, adding preceding and succeeding texts like an overlapping chunking method, give coherence to the evidence and help in better retrieval (users often use localizing words in questions that may not be exactly inside the pertinent evidence but could be nearby in the document). Surrounding texts clearly also help the LLM generate more consistent and relevant responses; (ii) Systematically, the page title (TTL) was found to be the single most helpful context to be added to the evidences ($0.483$ P@1 for TTL vs. $0.528$ for ALL, and $0.477$ Answer Relevance for TTL vs. $0.529$ for ALL), with each of the other three individual add-ons brings in substantial improvements over NONE (ALL ¿ TTL ¿ AFT ¿ BEF ¿ HDR ¿ NONE for P@1 and ALL ¿ TTL ¿ BEF ¿ HDR ¿ AFT ¿ NONE for answer relevance). No feature hurts performance. It is notable that succeeding content also helps, as common coreferencing methods only contextualize using preceding text. (iii) These improvements are systematic with respect to question complexity (simple/complex), answer source (passage/list/table), and language (English/German), demonstrating why contextualization is a worthwhile operation across diverse question types. An interesting thing to note here that the biggest gains in retrieval ($0.220$ gain in P@1 of ALL over NONE) and answering ($0.179$ gain in answer relevance of ALL over NONE) performance are achieved for answer-in-list questions – a very common QA scenario for enterprise wikis where one asks for meeting notes, attendees and action items – all often stored as lists.
@@
Counterfactual attribution is effective. While counterfactual evidence has been proven to be a tool for discovering causal factors (Pearl and Mackenzie, 2020), our particular implementation still needed proof of efficacy. Results in Table 3 (Column $4$) show that we consistently reach accuracies close to $80\%$, definitely an acceptable number in terms of user trust. The numbers in parentheses in Column $1$ show over how many questions the respective averages were computed in Columns $2-4$. For questions where the gold document/evidence was not retrieved in the top-$10$, there is no chance that the attribution module could spot the correct evidence. Hence these are removed from this evaluation and we do not see round numbers like $600/200/100$ as in Table 2. Looking at individual data slices across rows, it is satisfying to observe that the high performance was not due to skewed success on easier slices: improvements over standard or naive attribution are quite systematic over almost all criteria (varying only within a narrow span of $77-82\%$). In the naive mode (also a baseline in (Cohen-Wang et al., 2024)), we compute cosine similarities between the answer and the evidences, and then use softmax scores over this distribution as the attribution scores for the corresponding evidences. The top scoring evidence is used to compute accuracy (Sec. 7.1). Within the counterfactual approaches, the one with clustering generally does better (with clusters best in $5/8$ cases, without clusters in $2/8$: marked in bold in Table 3). Understandably, complex questions and cases where answers come from tables are slightly more difficult cases to solve (relatively lower accuracies of $\simeq 77\%$). Performance for German was slightly better than that for English.
@@ 7.4. In-depth RAG analysis
    [Table linearization] In Rows $10-12$, verbalization of table records indeed makes table contents more retrievable by a hybrid retriever (see below) as well as more digestible by an LLM, compared to natural linearization alternatives (the piped mode can be found in (Wang et al., 2024; Li et al., 2023a); markdown simply refers to the table in markdown format [^5] https://www.markdownguide.org/extended-syntax/#tables; HTML refers to the raw HTML of the table; plaintext is the remaining text of the table after stripping all HTML tags and attributes). But we also prefer verbalizations in the RAGonite interface owing to its easy scrutability by an end-user, over the HTML content of the table.
@@
    [Generation LLM] In Rows $16-18$, we find that using GPT-4o led to superior answer quality than Llama3.1 (we used the 8 billion parameter model [^6] https://huggingface.co/meta-llama/Llama-3.1-8B). But unfortunately the presence of large lists and table evidences in Confluence frequently led to the context window being exceeded for Llama, so its actual performance would be slightly better. The same LLM was used for both completed question and final answer generation. The P@1 is also lower for Llama ($0.480$ vs. $0.528$ for GPT-4o) because the retrieval in this case is carried out via the intent-explicit question generated by Llama, this is slightly inferior than the corresponding completions from GPT-4o. As requested in the prompt, the out-of-scope (OOS) message “The desired information cannot be found in the retrieved pool of evidence.” was triggered $199/600$ times ($33.2\%$) for the default RAGonite configuration with GPT-4o. Computing that the Precision@10 was $60.7\%$, an message would be expected $39.3\%$ of the times. Since this is very close to the actual measured rate of the OOS message, we can say that GPT-4o showed high instruction-following capabilities and did not use its parametric knowledge to generate its answers (undesirable in RAG).
@@
    [Embedding model] In Rows $19-21$, we observe that using bge-m3 embeddings for encoding text led to marginally better retrieval (P@1$=0.528$) for our data than the recent text-embedding-3-small model from OpenAI (P@1$=0.525$). While using the OpenAI model led to slightly better answer quality, better retrieval is the primary purpose of embeddings, leading to our choice of BGE as default. Moreover, BGE is used via local deployment, that we find preferable to an API for text-embedding-3-small because of latency and cost issues.
@@
    [Evidence ranking] In Rows $22-24$, hybrid search was found to be superior than lexical or dense search built inside ChromaDB. We used 10 evidences each from lexical and dense search to be fused for the hybrid variant. Hybrid search entails a re-ranking of evidences, as mentioned next.
@@
    [Evidence reranking] In Rows $25-27$, reranking evidences in the common pool of lexical and dense retrieval via the multilingual BGE re-ranking model to create the hybrid ranked list, was found to be better than a simple reciprocal rank fusion (RRF) approach. In RRF, the final list is created by ordering via a new evidence score that is the sum of the reciprocals of the ranks of the evidence in the lexical and dense rankings (a parameter $k=60$ is added to the original rank (Cormack et al., 2009) before computing the reciprocal). Note that RRF is always applied by default on the lexical and dense results to create the hybrid list: when the BGE reranker is applied, it works over the baseline list with RRF.
@@ 8. Conclusion and future work
RAGonite is an explainable RAG pipeline, that makes every intermediate step open to scrutiny by an end user. We propose flexible strategies for adding context to heterogeneous evidence, and causally explain LLM answers using counterfactual attribution. Future work involves fine-tuning specific RAG components for improving performance with smaller models.

### v3

@@ 6. The ConfQuestions Benchmark
Need for a new benchmark. While there are many public QA benchmarks (Roy and Anand, 2022), there were none that fit all our desiderata: one that uses heterogeneous pages in enterprise wikis, is suitable for ConvQA with RAG, and contains complex questions in more than one language. So we created our own benchmark ConfQuestions as follows. We crawled $10$ Spaces under https://openxt.atlassian.net/wiki/spaces, yielding $215$ public Confluence pages from the developer Atlassian. These contain many documents that are very close to each other content-wise (like test reports for software versions $6-9$, organizational policy over several years, and several meeting notes on a topic), yet each page contains unique information: this is very suitable for evaluating accurate retrieval and generation models. The pages contain a mix of concepts like software compatibility, build configurations, planning, along with mentions of people, dates and software, manifested via heterogeneous elements like tables, lists, and passages. These documents thus nicely simulate an enterprise Wiki space, our focus here.
Benchmark construction. The authors went through every page to judge the feasibility of generating NL conversations over them. Eventually, $50$ conversations were generated: $80\%$ with 5 turns, $20\%$ longer ones with $10$ turns, making context handling tougher, so $300$ questions in all. $20\%$ conversations also span two URLs instead of one, containing slight topic shifts. The annotators provided questions, intent-explicit versions, answer source (passage/list/table), complexity type (simple/complex), and ground truth answers along with source URLs. The questions were originally generated in English (EN). Native German speakers then translated both conversational and completed questions by hand to German (DE). We took special care to make the benchmark challenging from two aspects: (i) $50\%$ the questions are complex in the sense that they either require joining multiple evidences or involve aggregations, negations, or comparisons; and (ii) the answers to these questions are spread equally across passages, lists, and tables. So to be able to perform well overall on ConfQuestions, a RAG systems needs to be able to handle complex questions over heterogeneous elements in multiple languages. Benchmark statistics are in Table 1.
@@ 7. Results and Insights
Data $\downarrow$ / Contextualization $\rightarrow$ NONE +TTL +HDR +BEF +AFT +ALL NONE +TTL +HDR +BEF +AFT +ALL
@@ 7.1. Metrics
We use Precision@1 as our retrieval metric (since there is usually one gold URL), and Answer relevance as judged by GPT-4o (00 for non-relevant, $0.5$ for partially relevant, $1$ for relevant) against gold answers, as our metric for answer quality. Question-wise results are averaged over the full dataset. We use accuracy for evaluating attribution explanations as follows: we find the evidence that has the highest attribution score (for a cluster, the score of the cluster is considered to be the score of each evidence inside), and find its source URL. If this URL matches that of the gold answer, then accuracy is $1$, else 00. The average accuracy over all questions is then reported. This is interpretable as a percentage if multiplied by $100$.
For both Precision@1 and Attribution accuracy, we match the URL of the top-evidence against the answer URL in the benchmark. This is a proxy for ground truth: a perfect case would be to have gold labels at evidence-level, but this is very difficult to annotate for our benchmark, as gold answers are often composed of nuggets of information spread over the entire document.
@@ 7.2. Setup
We now run and evaluate RAGonite in various configurations over the $600$ English and German questions in ConfQuestions. The proposed default configuration uses: (i) all available context; (ii) all evidences as the corpus; (iii) LLM-generated intent-explicit questions; (iv) verbalization as the linearization technique for table records; (v) both row- and table-level indexing for table elements; (vi) GPT-4o as our LLM of choice; (vii) BGE as our text embedding model; (viii) hybrid search as our retrieval mode; and (ix) BGE+RRF reranking over the hybrid search results. We feed the reranked top-10 evidences ($k=10$) by the retriever into the LLM. Contextualization and attribution variants are discussed in Sec. 7.3: all other ablation analysis is postponed to Sec. 7.4.
Prompts. There are three prompts used in RAGonite: (i) for rephrasing a conversational question; (ii) for generating an answer from retrieved evidences; and (iii) for evaluating a generated answer. These are presented in Figs. 67, and 8, respectively (used for both English and German questions).
@@ 7.3. Key findings
Evidence contextualization improves performance. Contextualization results are in Table 2. We make the following observations: (i) Adding all the suggested contexts (ALL) consistently led to significantly improved performances for retrieval and generation. This means that enhancing evidences from documents with surrounding context is generally beneficial for RAG over collaborative wikis: content authors are often efficiency-oriented and do not repeat information inside passages, lists, or tables that are already mentioned in page titles or previous headings, for example. Similarly, adding preceding and succeeding texts, give coherence to the evidence and help in better retrieval (users often use localizing words in questions that may not lie exactly inside the pertinent evidence but could be nearby in the document). Surrounding texts clearly also help the LLM generate more consistent and relevant responses; (ii) Systematically, the page title (TTL) was found to be the single most helpful context to be added to the evidences ($0.483$ P@1 for TTL vs. $0.528$ for ALL, and $0.477$ Answer Relevance for TTL vs. $0.529$ for ALL), with each of the other three individual add-ons bringing in substantial improvements over NONE (ALL ¿ TTL ¿ AFT ¿ BEF ¿ HDR ¿ NONE for P@1 and ALL ¿ TTL ¿ BEF ¿ HDR ¿ AFT ¿ NONE for answer relevance). No feature hurts performance. It is notable that succeeding content also helps (a notable case is that of table footers), while common coreferencing methods only contextualize using preceding text; (iii) These improvements are systematic with respect to question complexity (simple/complex), answer source (passage/list/table), and language (English/German), demonstrating that contextualization is a worthwhile operation across diverse question types. We also note that the biggest gains in retrieval ($0.220$ gain in P@1 of ALL over NONE) and answering ($0.179$ gain in answer relevance of ALL over NONE) performance are achieved for answer-in-list questions – a very common QA scenario for enterprise wikis where one asks for meeting notes, attendees and action items – all often stored as lists.
@@
Counterfactual attribution is effective. While counterfactual evidence has been proven to be a tool for discovering causal factors (Pearl and Mackenzie, 2020), our particular implementation still needed proof of efficacy. Results in Table 3 (Column $4$) show that we consistently reach accuracies close to $80\%$, definitely an acceptable number in terms of user trust in RAGonite explanations. The numbers in parentheses in Column $1$ show over how many questions the respective averages were computed in Columns $2-4$. For questions where the gold document/evidence was not retrieved in the top-$10$, there is no chance that the attribution module could spot the correct evidence. Hence these are removed from this evaluation and we do not see round numbers like $600/200/100$ as in Table 2. Looking at individual data slices across rows, it is satisfying to observe that the high performance was not due to skewed success on easier slices: improvements over standard or naive attribution are quite systematic over almost all criteria (varying only within a narrow span of $77-82\%$). In the naive mode (also a baseline in (Cohen-Wang et al., 2024)), we compute cosine similarities between the answer and the evidences, and then use softmax scores over this distribution as the attribution scores for the corresponding evidences. The top scoring evidence is used to compute accuracy (Sec. 7.1). Within the counterfactual approaches, the one with clustering generally does better (with-clusters best in $5/8$ cases, without-clusters in $2/8$: marked in boldface in Table 3). Understandably, complex questions and answer-in-table questions are slightly more difficult cases to solve (relatively lower accuracies of $\simeq 77-78\%$). Performance for German was found to be slightly better than that for English.
@@ 7.4. In-depth RAG analysis
    [Table linearization] In Rows $10-12$, verbalization of table records indeed makes table contents more retrievable by a hybrid retriever (see below) as well as more digestible by an LLM, compared to natural linearization alternatives (the piped mode can be found in (Wang et al., 2024; Li et al., 2023a); markdown simply refers to the table in markdown format [^5] https://www.markdownguide.org/extended-syntax/#tables; HTML refers to the raw HTML of the table; plaintext is the remaining text of the table after stripping all HTML tags and attributes). But we also prefer verbalizations in the RAGonite interface owing to its scrutability by an end-user.
@@
    [Generation LLM] In Rows $16-18$, we find that using GPT-4o led to superior answer quality than Llama3.1-8B. Unfortunately the presence of large lists and table evidences in Confluence often led to the context window being exceeded for Llama. The same LLM was used for both completed question and final answer generation. The P@1 is also lower for Llama ($0.480$ vs. $0.528$ for GPT-4o) because the retrieval is carried out via the intent-explicit question generated by Llama (slightly inferior to corresponding completions from GPT-4o). As requested in the prompt, the out-of-scope (OOS) message “The desired information cannot be found in the retrieved pool of evidence.” was triggered $199/600$ times ($33.2\%$) for the default RAGonite configuration with GPT-4o. Computing the Precision@10 to be $60.7\%$, an OOS message would be expected $39.3\%$ of the times. Since this is very close to the actual measured rate of the OOS message, we can say that GPT-4o showed high instruction-following capabilities and did not use its parametric knowledge.
@@
    [Embedding model] In Rows $19-21$, we see that using bge-m3 embeddings for encoding text led to marginally better retrieval (P@1$=0.528$) than the recent text-embedding-3-small model from OpenAI (P@1$=0.525$). While using the OpenAI model led to slightly better answer quality, better retrieval is the primary purpose of embeddings, leading to our choice of BGE as default. Moreover, BGE is used via local deployment, that we find preferable to the OpenaI API due to latency and cost issues.
@@
    [Evidence ranking] In Rows $22-24$, hybrid search was found to be superior than lexical or dense search built inside ChromaDB. We used 10 evidences each from lexical and dense search.
@@
    [Evidence reranking] In Rows $25-27$, reranking evidences in the common pool of lexical and dense retrieval via the multilingual BGE reranking model+RRF to create the hybrid ranked list, was found to be better than using only RRF. In RRF, the final list is created by ordering via a new evidence score that is the sum of the reciprocals of the ranks of the evidence in the lexical and dense rankings (a parameter $k=60$ is added to the original rank (Cormack et al., 2009) before computing the reciprocal).
@@ 8. Conclusion and future work
RAGonite is an explainable RAG pipeline, that makes every intermediate step open to scrutiny by an end user. We propose flexible strategies for adding context to heterogeneous evidence, and causally explain LLM answers using counterfactual attribution. Future work will introduce cost and time-efficient agentic workflows.